# Spectral decomposition of field operators and causal measurement in quantum field theory

### v1

@@ 13 September 2024
We construct the spectral decomposition of field operators in bosonic quantum field theory as a limit of a strongly continuous family of POVM decompositions. The latter arise from integrals over families of bounded positive operators. Crucially, these operators have the same locality properties as the underlying field operators. We use the decompositions to construct families of quantum operations implementing measurements of the field observables. Again, the quantum operations have the same locality properties as the field operators. What is more, we show that these quantum operations do not lead to superluminal signaling and are possible measurements on quantum fields in the sense of Sorkin.
@@ 1 Introduction
In non-relativistic quantum mechanics, self-adjoint operators on Hilbert space play an important role in encoding physical observables that are used to describe measurement processes and their outcomes. The eigenvalues of the observable are the possible values labeling the outcomes of a measurement, while the eigenspaces determine the state after the measurement. That is, the spectral decomposition of observables plays an essential role in quantum measurement theory. In the special case of performing only a single measurement to obtain an expectation value of an observable and discarding the system afterwards one can get away without knowing the spectral decomposition of the observable explicitly. However, in the case of multiple measurements or a single measurement with post-selection this is no longer the case. Indeed, in modern quantum measurement theory, the basic concept is not that of an observable, but that of a quantum operation. This is a completely positive super-operator on the space of self-adjoint operators on Hilbert space. To construct the quantum operations corresponding to an observable requires precisely the spectral decomposition of the observable.
@@
Historically, observables in terms of self-adjoint operators have played a much less important role in describing measurement processes in quantum field theory. Instead, scattering theory has been of most interest, described through an asymptotic transition amplitude. This corresponds to a one-shot measurement where preparation and observation happen at asymptotically early and late times respectively. While this has been adequate for most applications in high energy physics, there is a growing need for a full-fledged measurement theory of quantum field theory that is able to describe processes involving multiple measurements that are moreover localized in both time and space. Rather than provide individual references on this subject we recommend to the reader an excellent recent survey article by Papageorgiou and Fraser [1].
@@
The most straightforward approach to a measurement theory of quantum field theory would be to follow the same path as in non-relativistic quantum mechanics. That is, use the spectral decomposition of the observables of interest and construct the corresponding quantum operations. The literature is surprisingly sparse on this topic. What is more, there is an important body of literature suggesting that this approach is problematic. We limit ourselves here to mention R. Sorkin’s seminal work on the subject, published in 1993 [2]. In it, Sorkin showed that a relatively generic class of projective measurements leads to superluminal signaling and is thus unphysical. A quantum operation constructed from projection operators of a spectral decomposition would presumably be affected by this.
@@
The simplest, but arguably most important class of localizable observables in QFT is given by field operators. By these we mean here those operators arising from the quantization of linear functions on phase space. Equivalently, these are the operators that can be written as the sum of a creation and corresponding annihilation operator. Using coherent states, properties of the Weyl quantization, and characteristic functions of measures, we motivate in Section 2 the sesquilinear form determining the spectral measure of field operators. In Section 3 we extend this sesquilinear form to a 1-parameter family and show that this gives rise to a 1-parameter family of positive-operator-valued measure (POVM) decompositions of field operators, satisfying continuity and composition properties, with the spectral measure as a limit. In Section 4 we construct quantum operations to measure the field operator. These come in two types. The discrete outcome type, based on the spectral decomposition, consists of partitioning the real line into subsets, with quantum operations testing whether the value of the observable lies in a given subset or not (Section 4.1). The continuous outcome type amounts to a full and precise measurement of the value of the observable (Section 4.2) and is based instead on the 1-parameter family of POVM decompositions, approximating the spectral decomposition. Locality and causality properties of the quantum operations are investigated in Section 5. In particular, we show that all quantum operations have the same locality properties as the corresponding field operator. Moreover, we show that measurements of the continuous outcome type do not lead to superluminal signaling, but satisfy what we call causal transparency. We conclude the paper with some discussion and an outlook in Section 6. An appendix contains part of the proof of Lemma 3.4.
@@ 2 Motivation
$\mathsf{K}_{\xi}(\phi)=\exp\left(\frac{1}{2}\{\xi,\phi\}\right),\quad\forall\xi\in L\,\forall\phi\in\hat{L}.$ (1)
@@ 3.1 Pointwise measure
Equation (20) is obtained by explicit calculation. It is valid as long as the denominators in the fractions are not zero. By the Cauchy-Schwarz inequality these are zero precisely if $\epsilon=\epsilon^{\prime}=0$ and moreover $\xi^{\prime}$ and $\xi$ are linearly dependent over $\mathbb{R}$, so that $\{\xi,\xi^{\prime}\}=\|\xi\|\|\xi^{\prime}\|$. If $\omega(\xi,\xi^{\prime})=0$, then $\{\xi,\xi^{\prime}\}=\{\xi^{\prime},\xi\}$ so that the right-hand side of equation (20) becomes symmetric under interchange of $(\xi,\epsilon,q)$ and $(\xi^{\prime},\epsilon^{\prime},q^{\prime})$. ∎
@@
That is, $B_{\xi}^{\epsilon}(q)$ is a positive multiple of a square of hermitian sesquilinear form and is thus positive. Positivity of $B_{\xi}^{0}(q)$ follows by taking the limit $\epsilon\to 0$ of $B_{\xi}^{\epsilon}(q)(\psi,\psi)$ for $\psi\in\mathcal{D}$ due to Lemma 3.2. ∎
@@ 3.2 Functional calculus
$\Pi_{\xi}^{\epsilon}[f]\coloneqq\int_{\-infty}^{\infty}\mathrm{d}qf(q)\Pi_{\xi}^{\epsilon}(q).$ (39)
@@ 4.2 Continuous outcomes
We now have at our disposal quantum operations that encode a precise measurement of the value of the observable $D_{\xi}$. For fixed $\epsilon>0$ the quantum operation that selects outcome $q\in\mathbb{R}$ is $A_{\xi}^{\epsilon}(q)$ as defined in expression (70). Since a single point $q$ has measure zero in $\mathbb{R}$ it generally does not make sense to use the quantum operation $A_{\xi}^{\epsilon}(q)$ directly. Instead, the relevant quantum operation is the weighted version $A_{\xi}^{\epsilon}[f]$ with weight function $f$, see expression (75). In particular, the non-selective quantum operation that corresponds to the measurement without reading any outcome is $A_{\xi}^{\epsilon}[\mathbf{1}]$. It is the requirement that $A_{\xi}^{\epsilon}[\mathbf{1}]$ be trace-preserving that fixes the numerical factor in definition (75). This has been carried over also to definition (70). Let $X\subseteq\mathbb{R}$ be Lebesgue measurable. Then, the selective quantum operation that tests whether the outcome lies in the subset $X$ is $A_{\xi}^{\epsilon}[\chi_{X}]$. On the other hand, the quantum operation encoding measurement of the expectation value is $A_{\xi}^{\epsilon}[\operatorname{id}]$.
@@
Another question is, how the quantum operations behave under change of the parameter $\epsilon$, and whether we can take a limit $\epsilon\to 0$. As to the first question, Lemma 4.6 provides a continuity result under a change of $\epsilon$. As to the second question, while it seems a limit $\lim_{\epsilon\to 0}A_{\xi}^{\epsilon}[f]$ as a quantum operation does not in general exist, this is not really necessary. What we need instead is that the limit of those expressions that yield probabilities or expectation values of measurements do exist. This is a much weaker requirement, weaker than the existence of the limit $\lim_{\epsilon\to 0}A_{\xi}^{\epsilon}[f](\sigma)$ for a fixed state $\sigma\in\mathcal{T}$. In this sense Lemma 4.8 is promising as it shows the existence of a limit under the trace.
@@ 5.1 Locality
Let $\xi\in L$ be localizable in a subset $S$ of a spacelike hypersurface $\Sigma$. Then, $\Pi_{\xi}[f]$ and $Q_{\xi}[X]$ are localizable is $S$. Also, for $\epsilon>0$, $\Pi_{\xi}^{\epsilon}(q)$, $\Pi_{\xi}^{\epsilon}[f]$, $A_{\xi}^{\epsilon}(q)$ and $A_{\xi}^{\epsilon}[f]$ are localizable in $S$.
Figure 1: Setup with two measurements $N$ and $M$. $N$ is non-selective, while $M$ is selective. By locality, the outcome of $M$ should not depend on whether or not $N$ is preformed.
@@ 5.2 Causal transparency
Figure 2: Setup with three measurements $N$, $I$ and $M$. $N$ and $I$ are non-selective, while $M$ is selective. Given locality, causal transparency of $I$ means that the outcome of $M$ should not depend on whether or not $N$ is preformed.
@@ 6 Discussion and Outlook
The support of the field operator is the union of the supports of the functions $f$ and $g$ on the hypersurface at $t$ in terms of our discussion of locality in Section 5.1. In particular, the elements of the phase space $L$ are in correspondence to pairs $(f,g)$ and allow for a rather straightforward interpretation in terms of initial data. On the other hand, it is often more convenient to use a smearing function $f$ in spacetime rather than space, which also allows to omit the operator $\hat{\dot{\phi}}(t,x)$. In this setting, variations of our continuous outcome measurement family have appeared previously in the literature in the context of weak measurements, see [9, 10] and references therein. What is more, I. Jubb has already argued that such measurements satisfy causal transparency [9].
R. Sorkin showed in a seminal work that a large class of projection valued measurements violate causal transparency [2]. Since Sorkin’s argument was quite generic, this has been taken to mean that any projection-valued measurement probably violates causal transparency. Since quantum operations corresponding to self-adjoint operators are normally constructed out of projectors arising from the spectral decomposition of the operator, this has been interpreted as an important obstacle in obtaining a reasonable theory of quantum measurement for quantum field theory [1]. In the context of the present work, this affects in particular the discrete outcome measurements discussed in Section 4.1. Indeed, the conclusion that they violate causal transparency has been reinforced in a recent more specific analysis of these types of measurements in Klein-Gordon theory by Albertini and Jubb [11]. This makes our proof of causal transparency for the continuous outcome measurements surprising, as one might think of these as arising from an (infinite) refinement of the discrete outcome measurements. In particular, our result suggests that a satisfactory theory of measurement, broadly analogous to that of non-relativistic quantum mechanics, but respecting locality and causality, can be constructed for quantum field theory after all.

### v2

@@ 13 September 2024
2 April 2025 (v2)
We construct the spectral decomposition of field operators in bosonic quantum field theory as a limit of a strongly continuous family of positive-operator-valued measure decompositions. The latter arise from integrals over families of bounded positive operators. Crucially, these operators have the same locality properties as the underlying field operators. We use the decompositions to construct families of quantum operations implementing measurements of the field observables. Again, the quantum operations have the same locality properties as the field operators. What is more, we show that these quantum operations do not lead to superluminal signaling and are possible measurements on quantum fields in the sense of Sorkin.
@@ 1 Introduction
In non-relativistic quantum mechanics, self-adjoint operators on Hilbert space play an important role in encoding physical observables that are used to describe measurement processes and their outcomes. The eigenvalues of the observable are the possible values labeling the outcomes of a measurement, while the eigenspaces determine the state after the measurement. That is, the spectral decomposition of observables plays an essential role in quantum measurement theory. In the special case of performing only a single measurement to obtain an expectation value of an observable and discarding the system afterwards one can get away without knowing the spectral decomposition of the observable explicitly. However, in the case of multiple measurements or a single measurement with post-selection this is no longer the case. Indeed, in modern quantum measurement theory, the basic concept is not that of an observable, but that of a quantum operation [1]. This is a completely positive super-operator on the space of self-adjoint operators on Hilbert space. To construct the quantum operations corresponding to an observable requires precisely the spectral decomposition of the observable.
@@
Historically, observables in terms of self-adjoint operators have played a much less important role in describing measurement processes in quantum field theory (QFT). Instead, scattering theory has been of most interest, described through an asymptotic transition amplitude. This corresponds to a one-shot measurement where preparation and observation happen at asymptotically early and late times respectively. While this has been adequate for most applications in high energy physics, there is a growing need for a full-fledged measurement theory of quantum field theory that is able to describe processes involving multiple measurements that are moreover localized in both time and space. Rather than provide individual references on this subject we recommend to the reader an excellent recent survey article by Papageorgiou and Fraser [2].
@@
The most straightforward approach to a measurement theory of quantum field theory would be to follow the same path as in non-relativistic quantum mechanics. That is, use the spectral decomposition of the observables of interest and construct the corresponding quantum operations. The literature is surprisingly sparse on this topic. What is more, there is an important body of literature suggesting that this approach is problematic. We limit ourselves here to mention Sorkin’s seminal work on the subject, published in 1993 [3]. In it, Sorkin showed that a relatively generic class of projective measurements leads to superluminal signaling and is thus unphysical. A quantum operation constructed from projection operators of a spectral decomposition would presumably be affected by this.
This difficulty has boosted efforts to look for alternative ways of describing measurement in QFT, particularly through an explicit modelling of the measurement apparatus as a quantum system in its own right which is subsequently measured in the conventional way, in the sense of von Neumann [4]. One would then aim to show that the interaction between system and apparatus satisfies locality and does not lead to superluminal signaling. Two approaches have reported success in this respect recently. One is based on particle-detector models [5], in the tradition of Unruh-DeWitt detectors [6]. The other one is based on modeling the apparatus through a quantum field [7], generalizing the work of Hellwig and Kraus [8]. (In the latter case discussion of Sorkin’s problem is in [9, 10].) Note that these approaches still need to recur to the conventional non-relativistic measurement scheme to extract the measured value from the apparatus after the system-apparatus interaction, potentially reintroducing vulnerabilities.
@@
The simplest, but arguably most important class of localizable observables in QFT is given by field operators. By these we mean here those operators arising from the quantization of linear functions on phase space. Equivalently, these are the operators that can be written as the sum of a creation and corresponding annihilation operator. Using coherent states, properties of the Weyl quantization, and characteristic functions of measures, we motivate in Section 2 the sesquilinear form determining the spectral measure of field operators. In Section 3 we extend this sesquilinear form to a 1-parameter family and show that this gives rise to a 1-parameter family of positive-operator-valued measure (POVM) decompositions of field operators, satisfying continuity and composition properties, with the spectral measure as a limit. In Section 4 we construct quantum operations to measure the field operator. These come in two types. The discrete outcome type, based on the spectral decomposition, consists of partitioning the real line into subsets, with quantum operations testing whether the value of the observable lies in a given subset or not (Section 4.1). The continuous outcome type amounts to a full and precise measurement of the value of the observable (Section 4.2) and is based instead on the 1-parameter family of POVM decompositions, approximating the spectral decomposition. Locality and causality properties of the quantum operations are investigated in Section 5. In particular, we show that all quantum operations have the same locality properties as the corresponding field operator. Moreover, we show that measurements of the continuous outcome type do not lead to superluminal signaling, but satisfy what we call causal transparency. We conclude the paper with some discussion and an outlook in Section 6. The appendix contains parts of the proofs of Lemmas 3.3 and 3.4.
@@ 2 Motivation
$\mathsf{K}_{\xi}(\phi)=\exp\left(\frac{1}{2}\{\xi,\phi\}\right),\quad\forall\xi\in L\,\forall\phi\in L.$ (1)
@@ 3.1 Pointwise measure
Equation (20) is obtained by explicit calculation, see Appendix A. It is valid as long as the denominators in the fractions are not zero. By the Cauchy-Schwarz inequality these are zero precisely if $\epsilon=\epsilon^{\prime}=0$ and moreover $\xi^{\prime}$ and $\xi$ are linearly dependent over $\mathbb{R}$, so that $\{\xi,\xi^{\prime}\}=\|\xi\|\|\xi^{\prime}\|$. If $\omega(\xi,\xi^{\prime})=0$, then $\{\xi,\xi^{\prime}\}=\{\xi^{\prime},\xi\}$ so that the right-hand side of equation (20) becomes symmetric under interchange of $(\xi,\epsilon,q)$ and $(\xi^{\prime},\epsilon^{\prime},q^{\prime})$. ∎
@@
That is, $B_{\xi}^{\epsilon}(q)$ is a positive multiple of a square of a hermitian sesquilinear form and is thus positive. Positivity of $B_{\xi}^{0}(q)$ follows by taking the limit $\epsilon\to 0$ of $B_{\xi}^{\epsilon}(q)(\psi,\psi)$ for $\psi\in\mathcal{D}$ due to Lemma 3.2. ∎
@@ 3.2 Functional calculus
$\Pi_{\xi}^{\epsilon}[f]\coloneqq\int_{-\infty}^{\infty}\mathrm{d}qf(q)\Pi_{\xi}^{\epsilon}(q).$ (39)
@@ 4.2 Continuous outcomes
We now have at our disposal quantum operations that encode a precise measurement of the value of the observable $D_{\xi}$. For fixed $\epsilon>0$ the quantum operation that selects outcome $q\in\mathbb{R}$ is $A_{\xi}^{\epsilon}(q)$ as defined in expression (70). Since a single point $q$ has measure zero in $\mathbb{R}$ it generally does not make sense to use the quantum operation $A_{\xi}^{\epsilon}(q)$ directly. Instead, the relevant quantum operation is the weighted version $A_{\xi}^{\epsilon}[f]$ with weight function $f$, see expression (75). In particular, the non-selective quantum operation that corresponds to the measurement without reading any outcome is $A_{\xi}^{\epsilon}[\mathbf{1}]$. It is the requirement that $A_{\xi}^{\epsilon}[\mathbf{1}]$ be trace-preserving that fixes the numerical factor in definition (75). This has been carried over also to definition (70). Let $X\subseteq\mathbb{R}$ be Lebesgue measurable. Then, the selective quantum operation that tests whether the outcome lies in the subset $X$ is $A_{\xi}^{\epsilon}[\chi_{X}]$. In particular, $X\mapsto A_{\xi}^{\epsilon}[\chi_{X}]$ defines a quantum instrument [17]. On the other hand, the quantum operation encoding measurement of the expectation value is $A_{\xi}^{\epsilon}[\operatorname{id}]$.
@@
Another question is, how the quantum operations behave under change of the parameter $\epsilon$, and whether we can take a limit $\epsilon\to 0$. As to the first question, Lemma 4.6 provides a continuity result under a change of $\epsilon$. As to the second question, while it seems a limit $\lim_{\epsilon\to 0}A_{\xi}^{\epsilon}[f]$ as a quantum operation does not in general exist, this is not really necessary. What we need instead is that the limit of those expressions that yield probabilities or expectation values of measurements do exist. This is a much weaker requirement, weaker even than the existence of the limit $\lim_{\epsilon\to 0}A_{\xi}^{\epsilon}[f](\sigma)$ for a fixed state $\sigma\in\mathcal{T}$. In this sense Lemma 4.8 is promising as it shows the existence of a limit under the trace.
@@ 5.1 Locality
Let $\xi\in L$ be localizable in a subset $S$ of a spacelike hypersurface $\Sigma$. Then, $\Pi_{\xi}[f]$ and $Q_{\xi}[X]$ are localizable in $S$. Also, for $\epsilon>0$, $\Pi_{\xi}^{\epsilon}(q)$, $\Pi_{\xi}^{\epsilon}[f]$, $A_{\xi}^{\epsilon}(q)$ and $A_{\xi}^{\epsilon}[f]$ are localizable in $S$.
Figure 1: Setup with two measurements $N$ and $M$. $N$ is non-selective, while $M$ is selective. By locality, the outcome of $M$ should not depend on whether or not $N$ is performed.
@@ 5.2 Causal transparency
Figure 2: Setup with three measurements $N$, $I$ and $M$. $N$ and $I$ are non-selective, while $M$ is selective. Given locality, causal transparency of $I$ means that the outcome of $M$ should not depend on whether or not $N$ is performed.
@@ 6 Discussion and Outlook
The support of the field operator is the union of the supports of the functions $f$ and $g$ on the hypersurface at $t$ in terms of our discussion of locality in Section 5.1. In particular, the elements of the phase space $L$ are in correspondence to pairs $(f,g)$ and allow for a rather straightforward interpretation in terms of initial data. On the other hand, it is often more convenient to use a smearing function $f$ in spacetime rather than space, which also allows to omit the operator $\hat{\dot{\phi}}(t,x)$. In this setting, variations of our continuous outcome measurement family have appeared previously in the literature in the context of weak measurements, see [18, 19] and references therein. What is more, Jubb has already argued that such measurements satisfy causal transparency [18].
Sorkin showed in a seminal work that a large class of projection valued measurements violate causal transparency [3]. Since Sorkin’s argument was quite generic, this has been taken to mean that any projection-valued measurement probably violates causal transparency. Since quantum operations corresponding to self-adjoint operators are normally constructed out of projectors arising from the spectral decomposition of the operator, this has been interpreted as an important obstacle in obtaining a reasonable theory of quantum measurement for quantum field theory [2]. In the context of the present work, this affects in particular the discrete outcome measurements discussed in Section 4.1. Indeed, the conclusion that they violate causal transparency has been reinforced in a recent more specific analysis of these types of measurements in Klein-Gordon theory by Albertini and Jubb [20]. This makes our proof of causal transparency for the continuous outcome measurements surprising, as one might think of these as arising from an (infinite) refinement of the discrete outcome measurements. In particular, our result suggests that a satisfactory theory of measurement, broadly analogous to that of non-relativistic quantum mechanics, but respecting locality and causality, can be constructed for quantum field theory after all.
@@ Acknowledgments
Statements and Declarations
No datasets have been analyzed or generated in this work. The author has no competing interests to declare that are relevant to the content of this article.
@@ Appendix A Appendix
of equation (20) of Lemma 3.3. We use the representation of the sesquilinear form given by equation (27) and the identity $D_{\xi}(\beta,\gamma)=\frac{\mathrm{i}}{2}(\{\beta,\xi\}-\{\xi,\gamma\})$ which follows from relations (4), (5) and (9).
$(B_{\xi^{\prime}}^{\epsilon^{\prime}}(q^{\prime})\star B_{\xi}^{\epsilon}(q))(\mathsf{K}_{\gamma},\mathsf{K}_{\beta})=\int_{\hat{L}}\mathrm{d}\nu(\phi)\,B_{\xi^{\prime}}^{\epsilon^{\prime}}(q^{\prime})(\mathsf{K}_{\gamma},\mathsf{K}_{\phi})B_{\xi}^{\epsilon}(q)(\mathsf{K}_{\phi},\mathsf{K}_{\beta})\\ =\frac{1}{\pi^{2}}\int_{\hat{L}}\mathrm{d}\nu(\phi)\int_{-\infty}^{\infty}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\,e^{-(\epsilon^{2}+\|\xi\|^{2})t^{2}-(\epsilon^{\prime 2}+\|\xi^{\prime}\|^{2})t^{\prime 2}}\langle\mathsf{K}_{\gamma},\mathsf{K}_{\phi}\rangle\,\langle\mathsf{K}_{\phi},\mathsf{K}_{\beta}\rangle\\ \exp\left(2\mathrm{i}t(D_{\xi}(\beta,\phi)-q)+2\mathrm{i}t^{\prime}(D_{\xi^{\prime}}(\phi,\gamma)-q^{\prime})\right)\\ =\frac{1}{\pi^{2}}\int_{-\infty}^{\infty}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\,e^{-(\epsilon^{2}+\|\xi\|^{2})t^{2}-(\epsilon^{\prime 2}+\|\xi^{\prime}\|^{2})t^{\prime 2}}\int_{\hat{L}}\mathrm{d}\nu(\phi)\langle\mathsf{K}_{\gamma},\mathsf{K}_{\phi}\rangle\,\langle\mathsf{K}_{\phi},\mathsf{K}_{\beta}\rangle\\ \exp\left(\frac{1}{2}\{2t\xi,\phi\}-\frac{1}{2}\{\phi,2t^{\prime}\xi^{\prime}\}+2\mathrm{i}t(D_{\xi}(\beta,0)-q)+2\mathrm{i}t^{\prime}(D_{\xi^{\prime}}(0,\gamma)-q^{\prime})\right)\\ =\frac{1}{\pi^{2}}\int_{-\infty}^{\infty}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\,e^{-(\epsilon^{2}+\|\xi\|^{2})t^{2}-(\epsilon^{\prime 2}+\|\xi^{\prime}\|^{2})t^{\prime 2}}\exp\left(2\mathrm{i}t(D_{\xi}(\beta,0)-q)+2\mathrm{i}t^{\prime}(D_{\xi^{\prime}}(0,\gamma)-q^{\prime})\right)\\ \int_{\hat{L}}\mathrm{d}\nu(\phi)\langle\mathsf{K}_{\gamma-2t^{\prime}\xi^{\prime}},\mathsf{K}_{\phi}\rangle\,\langle\mathsf{K}_{\phi},\mathsf{K}_{\beta+2t\xi}\rangle\\ =\frac{1}{\pi^{2}}\int_{-\infty}^{\infty}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\,e^{-(\epsilon^{2}+\|\xi\|^{2})t^{2}-(\epsilon^{\prime 2}+\|\xi^{\prime}\|^{2})t^{\prime 2}}\exp\left(2\mathrm{i}t(D_{\xi}(\beta,0)-q)+2\mathrm{i}t^{\prime}(D_{\xi^{\prime}}(0,\gamma)-q^{\prime})\right)\\ \langle\mathsf{K}_{\gamma-2t^{\prime}\xi^{\prime}},\mathsf{K}_{\beta+2t\xi}\rangle\\ =\frac{1}{\pi^{2}}\int_{-\infty}^{\infty}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\,e^{-(\epsilon^{2}+\|\xi\|^{2})t^{2}-(\epsilon^{\prime 2}+\|\xi^{\prime}\|^{2})t^{\prime 2}}\exp\left(2\mathrm{i}t(D_{\xi}(\beta,0)-q)+2\mathrm{i}t^{\prime}(D_{\xi^{\prime}}(0,\gamma)-q^{\prime})\right)\\ \langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\exp\left(\frac{1}{2}\{2t\xi,\gamma\}-\frac{1}{2}\{\beta,2t^{\prime}\xi^{\prime}\}-2\{\xi,\xi^{\prime}\}tt^{\prime}\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\pi^{2}}\int_{-\infty}^{\infty}\mathrm{d}t\int_{-\infty}^{\infty}\mathrm{d}t^{\prime}\,e^{-(\epsilon^{2}+\|\xi\|^{2})t^{2}-(\epsilon^{\prime 2}+\|\xi^{\prime}\|^{2})t^{\prime 2}-2\{\xi,\xi^{\prime}\}tt^{\prime}}\\ \exp\left(2\mathrm{i}t(D_{\xi}(\beta,\gamma)-q)+2\mathrm{i}t^{\prime}(D_{\xi^{\prime}}(\beta,\gamma)-q^{\prime})\right)\\ =\langle\mathsf{K}_{\gamma},\mathsf{K}_{\beta}\rangle\frac{1}{\pi\sqrt{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}}\\ \exp\left(\frac{1}{(\|\xi\|^{2}+\epsilon^{2})(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})-\{\xi,\xi^{\prime}\}^{2}}\left(2\{\xi,\xi^{\prime}\}\left(D_{\xi}(\beta,\gamma)-q\right)\left(D_{\xi^{\prime}}(\beta,\gamma)-q^{\prime}\right)\right.\right.\\ \left.\left.-(\|\xi^{\prime}\|^{2}+\epsilon^{\prime 2})\left(D_{\xi}(\beta,\gamma)-q\right)^{2}-(\|\xi\|^{2}+\epsilon^{2})\left(D_{\xi^{\prime}}(\beta,\gamma)-q^{\prime}\right)^{2}\right)\right).$ (93)
Proof.